OpenStudy (anonymous):

How do I write this equation in standard form? y = 2 - 4x

OpenStudy (loser66):

It is.

OpenStudy (anonymous):

standard form: Ax+By=C

OpenStudy (anonymous):

Thanks Taylor, but how do I convert it that way?

OpenStudy (anonymous):

ot is currently in slope-intercept form, so what do you do first

OpenStudy (anonymous):

I have already attempted it and it may be wrong, but is this the final answer? -4x + y = -2

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

yes so you know how to do it?

OpenStudy (anonymous):

Wait, it's this 4x - y = -2

OpenStudy (anonymous):

no, it is 4x-y=2

OpenStudy (anonymous):

no the two doesn't move ti stays negative

OpenStudy (anonymous):

My homework doesn't have an answer choice for 4x-y=2

OpenStudy (anonymous):

it was never negative in the first place

OpenStudy (anonymous):

ohhh unless you moved it, then it would be -2

OpenStudy (anonymous):

yeah I think the 2 is negative.

OpenStudy (anonymous):

oh loli didn't see he top quesiton your right

OpenStudy (anonymous):

It is, and it should be correct

OpenStudy (anonymous):

I will let you know if it is. Thank you both.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

You are very welcome!

OpenStudy (triciaal):

y = 2 - 4x 4x + y = 2